OpenStudy (anonymous):

Need confirmation on my explanation of the solution. I think i understand but since the book doesn't clearly explain it i want to know if my reasoning is correct. (Question will follow shortly).

OpenStudy (anonymous):

Here's the question: a) Find the solution of the given initial value problem in explicit form b) plot a graph of the solution c) determine (at least approximately) the interval in which the solution is defined. \[\frac{ \delta r( \theta ) }{ \delta \theta } =\frac{ r^2 (\theta ) }{ \theta }\] r(1)=2 now the general solution is \[\frac{ 2 }{ 1-2\ln( \theta ) } \in 0 < \theta < \sqrt{e}\] now there's two things i would like to confirm. Let's call the Q and W Q) The function above does not actually end after Sqrt(e), however it has no positive value after that point. Since the question originally asked for a r^2 i guess negative solutions don't count and that's why the solutions end at sqrt(e), is that correct? W) when \[\int\limits_{\infty}^{-\infty} \frac{ 1 }{ \theta } \delta \theta \] normally you'd get \[\ln \left| \theta \right| \] however in this case \[\int\limits_{}^{} \frac{ 1 }{ r^2 } dr = \int\limits_{}^{} \frac{ 1 }{ \theta } d \theta \] you get simply \[\ln \theta \] is that because \[\frac{ -1 }{ r }=\ln \left| \theta \right| \] would not make sense?

OpenStudy (anonymous):

Actually the solution i posted is not the general solution but the initial value solution. but that doesn't really matter for the question.

terenzreignz (terenzreignz):

Might I try to solve it myself (the differential equation) ?

OpenStudy (anonymous):

Sure, go ahead.

OpenStudy (anonymous):

In case it isn't clear from context it's actually r^2 (as a function of theta) on the top function, not r^2 multiplied by theta.

OpenStudy (anonymous):

how do you get your equations large and colored btw? mine are always so small and hard to read sometimes.

terenzreignz (terenzreignz):

\large \Large \LARGE \huge \Huge these are varying degrees of font size, type them before the equations to adjust size. \color{colour you want}{coloured text} colors that I know work are white, black, yellow, blue, red, green, violet, orange... I don't really see why you'd need any more than those ^_^ now...

OpenStudy (anonymous):

btw you made a mistake at your first equation

terenzreignz (terenzreignz):

What's that?

OpenStudy (anonymous):

\[\large \frac{ dr }{ r^2 } = \frac{ d \theta }{ \theta } \]

terenzreignz (terenzreignz):

Oh yeah... crud. Sorry about that.

terenzreignz (terenzreignz):

See why you shouldn't consult me? I'm error prone XD

terenzreignz (terenzreignz):

\[\Large -\frac1r= \ln|\theta|+C \]

OpenStudy (anonymous):

wait did you just edit your post?

terenzreignz (terenzreignz):

I didn't. Should I have?

terenzreignz (terenzreignz):

I just corrected it by posting under it.

terenzreignz (terenzreignz):

Posts cannot be edited, btw, only deleted.

OpenStudy (anonymous):

oh ok, i thought i went crazy when all of a sudden the wrong equation was replaced by a new one i guess.

OpenStudy (anonymous):

anyway yes terenz that is what logically i would say as well, however it actually incorrect.

OpenStudy (anonymous):

The ln does not have absolute signs in this case. Let me try if i can prove it somehow.

terenzreignz (terenzreignz):

What do you mean it doesn't have absolute signs?

terenzreignz (terenzreignz):

Why shouldn't it?

OpenStudy (anonymous):

well

OpenStudy (anonymous):

that's what i was trying to figure out as well

OpenStudy (anonymous):

In fact i think for the given domain it doesn't even matter.

terenzreignz (terenzreignz):

What do you mean it doesn't end at \(\large \sqrt e \)?

OpenStudy (anonymous):

Well, if you draw the graph while ignoring the limits given, the graph just continues (after the asymptotic behavior) on the negative axis

terenzreignz (terenzreignz):

Yes it does, so it's fair to say the function simply isn't defined at \(\large \theta = \sqrt e\) Should it be?

OpenStudy (anonymous):

Yes it should.

terenzreignz (terenzreignz):

It should? What is the apparent value of r when \(\large \theta = \sqrt e\)?

OpenStudy (anonymous):

Oh in fact no, i misinterpret your question. It's undefined at that point because division by 0

terenzreignz (terenzreignz):

Okay, so it's undefined at that point, is that a problem? :)

OpenStudy (anonymous):

It's a bit hard to see from these graphs but both the ln with and without absolute marks actually completely overlap the initial value problem of the differential equation.

OpenStudy (anonymous):

No it's not a problem, i just wanted to confirm if my reasoning is true. The graph of the solution ends after sqrt{e} is that because r^2 doesnt allow negative solutions?

terenzreignz (terenzreignz):

No, it doesn't have to end with \(\sqrt e\) r^2 concerns the derivatice, and if that's the case, it only means the function r is increasing at all times (which it is, check the graph)

terenzreignz (terenzreignz):

derivative^

OpenStudy (anonymous):

but the initial value solution of the equation actually does stop after sqrt(e) why exactly is that?

terenzreignz (terenzreignz):

Where does it say that it stops at sqrt(e)?

OpenStudy (anonymous):

The graph shows so

terenzreignz (terenzreignz):

It's not that it stops, it's more like it has an asymptote at \(\theta = \sqrt e\) But it is also defined for values \(\theta > \sqrt 3\)

terenzreignz (terenzreignz):

As the graph also clearly indicates.

OpenStudy (anonymous):

Which one?

terenzreignz (terenzreignz):

Your wolfram graph.

OpenStudy (anonymous):

which of the ones? the one is from the solution, the other one is from the differential equation. As you can see the differential equation does not have negatives, while the solution does (if you don't exclude values greater than sqrt(e)

OpenStudy (anonymous):

Here's all those 3 graphs plotted on my calculator http://imageshack.us/photo/my-images/35/7cw.JPG/

terenzreignz (terenzreignz):

The differential equation has no negatives for all \(\theta > 0\) yes?

OpenStudy (anonymous):

It doesn't even exsist for \[0 > \theta > \sqrt{e} \]

terenzreignz (terenzreignz):

We can ignore the negative values of \(\theta\) But for \(\theta > \sqrt e\), it does in fact exist. The wolfram graph shows it....

terenzreignz (terenzreignz):

Why is that so hard to believe?

OpenStudy (anonymous):

Then why does the solution stop after that point?

terenzreignz (terenzreignz):

Who said the solution stopped after that point?

OpenStudy (anonymous):

well for one the textbook, for two my calculator. also, i think when you refer to wolfram, you are not refering to the graph of the differential.

terenzreignz (terenzreignz):

The textbook explicitly says it is not defined for values greater than \(\sqrt e\) ?

terenzreignz (terenzreignz):

Or is it just the graph?

OpenStudy (anonymous):

the textbook states it as i said, it's only defined between 0 and sqrt(e) i just wanted to confirm if my logic is not flawed. I didn't want to know if the solution is wrong or not, i wanted to know the REASON behind the solution is correct or not.

terenzreignz (terenzreignz):

Perhaps it was simply restricted to that domain?

OpenStudy (anonymous):

it's the interval at which the solution is defined. that's what the textbook says.

terenzreignz (terenzreignz):

Then I'm at a loss. I still see no reason why it can't be defined for \(\theta > \sqrt e\) however.

OpenStudy (anonymous):

because at that point you divide by zero and after that point the graph will continue in the negative. Which is NOT what the differential equation graph does.

terenzreignz (terenzreignz):

You do not divide by zero when \(\theta > \sqrt e\) only when \(\theta = \sqrt e\)

OpenStudy (anonymous):

yes that's what i meant, at that point (sqrt(e)) you divide by zero and after that point the solutions will all be negative (asymptotic towards 0).

terenzreignz (terenzreignz):

And what's wrong when the solutions are negative?

OpenStudy (anonymous):

The differential equation only goes UP from that point, and at Sqrt(e) the slope is infinite, so it will never pass that point. Hence, they will never reach negative values. Nor anything to the right of that value.

OpenStudy (anonymous):

If however we would have picked another initial value then negatives would have been possible.

terenzreignz (terenzreignz):

Okay... assuming that, what's wrong now? You seem to have everything under control...

OpenStudy (anonymous):

I guess so, although it doesn't feel like i understand. Anyway thanks for your help. Talking about it did make it somewhat clearer i guess.

terenzreignz (terenzreignz):

Though, at sqrt(e), the slope is infinite, still, at the right-side of the sqrt(e) So what's wrong with having the slope decrease from then on?

OpenStudy (anonymous):

I will leave the question open for a bit in case people want to add to it to clear it up a bit but it has no hurry, i'll move on to other problems from the book now.

OpenStudy (anonymous):

We have an initial value problem, the graph that goes trough the initial value will not ever reach that point. So we can disregard all the points that will not be reached by the initial value?